8/(y+4)-6/(y-4)=8/(y^2-16) hey this problem has been giving me trouble if you could please show me the steps to. Solve so I understand better thank you.
Answer by stanbon(75887)

8/(y+4)-6/(y-4)=8/(y^2-16)
----
Multiply thru by (y+4)(y-4) to get:
8(y-4) - 6(y+4) = 8
----
8y-32 -6y-24 = 8
----
2y - 56 = 8
2y = 64
y = 32
===========
Cheers,
Stan H.
============